# Frequency-Domain Characterization of Load Demand from Electrified Highways

### v1

@@ V Periodograms of DWPT Load Demands
Periodogram under S4). To simulate S4), EV speeds $v_{n}$ were drawn independently from a Gaussian PDF with $\mu_{v}=24.56$ m/s ($55$ mph) and $\sigma_{v}=1.23$ m/s ($2.75$ mph), while EV timings $t_{n}$ were iid and uniform in $0,D/v_{n})$. Figure 4 compares the periodograms obtained under S4) to S2) and S3) for $N=192$ EVs. Unequal speeds cause spectrum spreading under S4). As asserted by Lemma 4, the bandwidth of the bells increases with $m$. The magnitude of the peaks is smaller under S4) rather than S2), as expected. The test also corroborates that the THCs for S2) and S4) are equal on average.
VI Tests Using the SUMO Simulator
@@
We simulated an $L=4$ km ($2.5$ mile) ER segment with two lanes; the rightmost one is the DWPT lane. Vehicles were allowed to change lanes. All vehicles driving in the DWPT lane were considered to be EVs of the same type (class-8 trucks). The DWPT parameters $(\ell_{T},\ell_{R},d)$ matched the INDOT testbed specifications. EV power densities were drawn independently from a Gaussian distribution with mean $\mu_{a}=94.45$ kW/m and standard deviation $\sigma_{a}=5.55$ kW/m. For SUMO parameters, we set the maximum vehicle speed to be $36$ m/s ($80$ mph), while vehicle acceleration was constrained to lie within $4.5,2.6]~\text{m/s}^{2}$. Successive vehicles were spaced by more than $2.5$ m and $1$ sec. Traffic followed the Krauss car-following model [24]. According to this model, the driver of vehicle $n$ has a desired speed $v_{n}^{d}$, but may have to slow down depending on the speed of the leading vehicle.
@@
Figure 5: Periodograms computed using 1-min DWPT loads simulated using SUMO under light traffic conditions. The periodograms contain bells centered around harmonics of $f_{0}=5.38$ Hz associated with mean speed $\mu_{v}=24.56$ m/s ($55$ mph) and the INDOT testbed segment spacing. Bells decay in magnitude and spread in bandwidth with increasing $m$, as asserted by Lemma 4. Periodograms across realizations are similar in shape.
@@
Free-flowing traffic conditions. With our first test using SUMO, we wanted to test whether S4) can explain the spectrum of the DWPT load under light traffic. To this end, we simulated a mean arrival rate of $\lambda_{a}=0.21$ EV/s. The desired EV speeds $v_{n}^{d}$ were drawn from a truncated Gaussian distribution with mean $\mu_{v}=24.56$ m/s, standard deviation $\sigma_{v}=2.45$ m/s ($5.5$ mph), while the lower and upper cutoff speeds were set to $20.11$ m/s ($45$ mph) and $29.05$ m/s ($65$ mph), respectively. We simulated traffic for over 40 min, and then randomly selected 1-min intervals as $\hat{p}(t)$. The total number of EVs driving on the ER segment remained relatively constant at $N=49$ EVs across all 1-min intervals; see Remark 2. Figure 5 depicts the periodograms obtained over 6 Monte Carlo runs. The DC component is $6.8$ MW, the power contained around the first harmonic is $613.70$ kW. THC in the range of 6–7% was observed for different $\hat{p}(t)$ segments. Overall, the key features of the spectra in Fig. 5 agree with the analysis under S4).
Figure 6: Variation of THC across time under low-speed, high-volume traffic simulated in SUMO. The speed affects the number of EVs and the EV spacing. Some speeds result in near-synchronization and high THC, as in S1), and other speeds result in lower THC, as in S4).
@@
Congested traffic conditions. With the next test, we wanted to see if there exist traffic conditions under which the DWPT demand exhibits higher THC due to EV synchronization as in S1). To this end, we studied the DWPT load on an ER with heavy traffic moving at a slower speed. Such conditions can arise due to an accident, adverse weather conditions, or construction activities. To simulate such traffic conditions in SUMO, the speed limit was $24.56$ m/s ($55$ mph) on the first half of the ER and lowered on the second part. Initial EV speeds were drawn from the truncated Gaussian distribution explained earlier. The mean arrival rate was set to $\lambda_{a}=1$ EV/s. At this rate, the traffic jam propagated from the second half backwards to the beginning of the ER. Eventually, all EVs were moving at roughly the slower speed limit of the second half. Inter-vehicle spacings were found to be approximately constant across vehicles and time.
@@
As explained under S1), EV loads at constant speed can synchronize only when EV spacings are integer multiples of the coil segment length $D$. Moreover, the distance gap between EVs increases with the speed limit [24]. Based on this understanding, we identified several speed limits for the second half of the ER, at which approximate synchronization occurs and leads to high THC. Figure 6 depicts how the THC of the total DWPT load demand evolves across time for different jam speeds. The THC was computed based on $1$-min segments of $p(t)$ with a 50% overlap. We observe that higher speeds result in fewer EVs, as expected based on Remark 2. We also note that the THC scales as $\text{THC}_{h}/\sqrt{N}$ for the bottom three speed limits, for which the inter-vehicle distance was not a multiple of $D$. On the contrary, for the top three speed limits, THC increases due to the resonance of traffic with the ER infrastructure. The maximum THC value we observed was $18\%$ for a particular $1$-minute segment when the jam moves at $7.81$ m/s ($17.47$ mph). Overall, this experiment validates the conjecture made in S1) that the DWPT load demand can exhibit high THC under heavy traffic conditions moving at specific speeds, especially at relatively lower speed limits. Also, the THC appears to be varying sinusoidally, which is a common phenomenon for frequency-modulated signals [20]. Analyzing the PSD under time-varying velocities and a detailed explanation of this sinusoidal variation in THC would be the topic of future study.
VII Conclusions
@@
This work puts forth statistical models for the total load to be consumed at the substation serving an ER. By postulating different distributions on EV timings, our analysis explains how the DWPT load spectrum varies under various traffic conditions. The worst-case scenario in terms of THC occurs when EVs move in a synchronized fashion with particular EV spacings. Simulations using SUMO demonstrate that near synchronization can occur under low-speed, high-volume traffic. When EVs move freely, the DWPT load exhibits a more favorable spectrum profile, as verified by simulating low-volume traffic in SUMO. Frequency content at non-DC frequencies scales with $\sqrt{N}$ or $N$ depending on EV synchronism. Platoon formations can yield higher THC. In general, the fundamental frequency depends on the spacing between coils and the average EV speed, so that low-speed traffic can result in harmonics below 2 Hz. Higher-order harmonics decrease in magnitude and spread across a wider bandwidth. Despite the simplicity of the models, our findings provide valuable insights to ER planners and grid operators on how the DWPT spectrum scales with the number of EVs. This work can be expanded in various directions: a) Model DWPT load when EVs move with time-varying speeds using tools from frequency-modulated signals; b) Study the effect of correlated EV speeds based on dynamical models of traffic flows; c) Analyze the effect of DWPT load consumed at multiple buses on grid frequency dynamics; and d) Develop solutions to smooth out DWPT load.
VIII Appendix
@@
where $a_{i}$ is defined as the effective rating of platoon $i$. Based on this decomposition, the DWPT load is now amenable to the analysis of Lemma 2 subject to two modifications. First, signal $p(t)$ is now the sum of $N/Q$ rather than $N$ stochastic components, each with iid and uniformly distributed timings $t_{i}$. Second, the effective rating $a_{i}$ of each stochastic component has a mean $Q\mu_{a}$ and its variance is increased to $Q\sigma_{a}^{2}$. The mean and autocorrelation of $p(t)$ can be obtained as in Lemma 2 by simply replacing $N$ by $N/Q$, $\mu_{a}$ by $Q\mu_{a}$, and $\sigma_{a}^{2}$ by $Q\sigma_{a}^{2}$. The autocorrelation function is

### v2

@@ V Periodograms of DWPT Load Demands
Periodogram under S4. To simulate S4, EV speeds $v_{n}$ were drawn independently from a Gaussian PDF with $\mu_{v}=24.56$ m/s ($55$ mph) and $\sigma_{v}=1.23$ m/s ($2.75$ mph), while EV timings $t_{n}$ were i.i.d. and uniform in $0,D/v_{n})$. Figure 5 compares the periodograms obtained under S4 to S2 and S3 for $N=192$ EVs. Unequal speeds cause spectrum spreading under S4. As asserted by Lemma 4, the bandwidth of the lobes increases with $m$. The magnitude of the peaks is smaller under S4 rather than S2, as expected. The test also corroborates that the THCs for S2 and S4 are equal on average. In theory, the first $8$ lobes should not overlap as the condition $\mu_{v}>1.15(2m+1)\sigma_{v}$ is satisfied. Some overlap is observed here due to windowing and finite-sample effects.
VI EVs Moving at Time-Varying Speeds
This section studies the spectrum of individual EVs moving at time-varying speeds.
VI-A Scenario S5: EV Moving at Time-Varying Speed
The instantaneous speed of EVn can be decomposed as
$v_{n}(t)=\bar{v}_{n}+\tilde{v}_{n}(t)$
where $\tilde{v}_{n}(t)$ is the deviation of the instantaneous speed from its time-averaged speed $\bar{v}_{n}$. Then, the position of EVn at time $t$ can be expressed as
$x_{n}(t)=\bar{v}_{n}t+\int_{0}^{t}\tilde{v}_{n}(\tau)d\tau.$ (23)
According to (3), the power drawn by EVn is a periodic function of $x_{n}$ and therefore is amenable to the FS expansion
$p_{n}(t)=\bar{p}_{n}(x_{n}(t))=a_{n}\sum_{m=-\infty}^{\infty}c_{m}e^{jm\omega_{s}x_{n}(t)},$ (24)
where $\omega_{s}=2\pi/D$ is the spatial frequency . Substituting (23) into (24) yields
$p_{n}(t)=a_{n}\sum_{m=-\infty}^{\infty}c_{m}e^{jm\omega_{s}\bar{v}_{n}t}\cdot e^{jm\omega_{s}\int_{0}^{t}\tilde{v}_{n}(\tau)d\tau}.$ (25)
For general aperiodic signals $\tilde{v}_{n}(t)$, it is difficult to characterize the spectrum of $p_{n}(t)$. However, if $\tilde{v}_{n}(t)$ is approximated as periodic with frequency $\omega_{n}=2\pi/T_{n}$, the second complex exponential in (25) is also amenable to an FS expansion as [26, Sec 3.3.2]:
$e^{jm\omega_{s}\int_{0}^{t}\tilde{v}_{n}(\tau)d\tau}=\sum_{k=-\infty}^{\infty}\beta_{k}(m)e^{jk\omega_{n}t},$ (26)
where $\beta_{k}(m)$ denotes the $k$-th FS coefficient corresponding to the $m$-th harmonic. Reference [26, Sec 3.3.2] computes these coefficients for $\tilde{v}_{n}(t)$ being a sinusoidal signal. Although there is no closed-form expression for $\beta_{k}(m)$ for general periodic $\tilde{v}_{n}(t)$, the coefficients $\beta_{k}(m)$ can be shown to decrease exponentially with $k$. Substituting (26) into (25) provides
$p_{n}(t)=a_{n}\sum_{m=-\infty}^{\infty}\sum_{k=-\infty}^{\infty}c_{m}\beta_{k}(m)e^{j(m\omega_{s}\bar{v}_{n}+k\omega_{n})t}.$ (27)
Therefore, the spectrum of $p_{n}(t)$ consists of peaks at frequencies $m\omega_{s}\bar{v}_{n}+k\omega_{n}$ for all $k$ and $m$. However, the peaks corresponding to larger $k$’s have low amplitude as the coefficients $\beta_{k}(m)$ decrease in $k$. This special case shows the technical difficulty in characterizing the spectrum of a single EV load, let alone an aggregation of them. However, the spectral spreading due to time-varying speeds can be observed numerically by computing spectra of DWPT loads derived from real-world traffic tests in Section VII.
VII Tests Using the SUMO Simulator
@@
We simulated an $L=4$ km ($2.5$ mile) ER segment with two lanes. The DWPT lane is the rightmost one. Vehicles were allowed to change lanes so they could overtake the proceeding vehicle if needed by moving temporarily to the non-DWPT leftmost lane. All vehicles driving in the DWPT lane were considered to be EVs of the same type (class-8 trucks). The DWPT parameters $(\ell_{T},\ell_{R},d)$ matched the INDOT testbed specifications. EV power densities were drawn independently from a Gaussian distribution with mean $\mu_{a}=94.45$ kW/m and standard deviation $\sigma_{a}=5.55$ kW/m. For SUMO parameters, we set the maximum vehicle speed to be $36$ m/s ($80$ mph), while vehicle acceleration was constrained to lie within $4.5,2.6]~\text{m/s}^{2}$. Successive vehicles were spaced by more than $2.5$ m and $1$ sec. Traffic followed the Krauss car-following model [16]. According to this model, the driver of vehicle $n$ has a desired speed $v_{n}^{d}$, but may have to slow down depending on the speed of the leading vehicle.
@@
Figure 6: Periodograms computed from 1-min DWPT loads simulated using SUMO under light-traffic conditions. The periodograms contain lobes centered around the harmonics of $f_{0}=5.38$ Hz associated with mean speed $\mu_{v}=24.56$ m/s ($55$ mph) and the INDOT testbed segment spacing. lobes decay in magnitude and spread in bandwidth with increasing $m$, as asserted by Lemma 4. Periodograms across realizations are similar in shape. lobes start overlapping beyond the third harmonic.
@@
Free-flowing traffic conditions. With our first test using SUMO, we wanted to test whether S4 can explain the spectrum of the DWPT load under light traffic. Additionally, we demonstrate the spreading of the spectra due to EVs moving with time-varying speeds as highlighted in S5. To this end, we simulated a mean arrival rate of $\lambda_{a}=0.21$ EV/s. The desired EV speeds $v_{n}^{d}$ were drawn from a truncated Gaussian distribution with mean $\mu_{v}=24.56$ m/s, standard deviation $\sigma_{v}=2.45$ m/s ($5.5$ mph), while the lower and upper cutoff speeds were set to $20.11$ m/s ($45$ mph) and $29.05$ m/s ($65$ mph), respectively. We simulated traffic for over 40 min, and then randomly selected 1-min intervals as $\hat{p}(t)$. The total number of EVs driving on the ER segment remained relatively constant at $N=49$ EVs with a maximum variation of $7\%$ across all 1-min intervals; see Remark 2. Figure 6 depicts the periodograms obtained over 6 Monte Carlo runs. The DC component is $6.8$ MW, the power contained around the first harmonic is $613.70$ kW. THC in the range of 6–7% was observed for different $\hat{p}(t)$ segments.
Because vehicles move under free-flow traffic conditions, their actual speeds are close to their respective desired speeds $v_{n}^{d}$. Therefore, for all 1-min realizations $\hat{p}(t)$, EV instantaneous speeds can roughly be described by the aforementioned Gaussian distribution. Because EV average speeds $\bar{v}_{n}$ are close to $v_{n}^{d}$, the spectra contains main lobes at frequencies $\mu_{v}/D$. Additional sidelobes appear as EVs move with time-varying speeds as postulated in S5. Overall, the key features of the spectra in Fig. 6 agree with the analysis under S4 and S5.
Figure 7: DWPT periodograms computed over two different window durations $T_{p}$. Longer windows enhance frequency resolution as the number of time samples increases.
The next test studies how the signal duration $T_{p}$ affects the periodogram $\hat{S}_{p}(\omega)$. As explained earlier, the periodogram is expected to converge to the PSD when the signal $\hat{p}(t)$ is of infinite duration. To that end, we compare the frequency content of the power drawn by EVs under free-flowing traffic conditions by evaluating the PSD of DWPT loads of different time resolution (See Fig. 7). We observe that increasing the signal duration leads to sharper peaks. This is because longer time duration enhances frequency resolution.
Figure 8: Spectrogram of DWPT loads when EVs move in free-flowing conditions. Observe that frequency content remains approximately constant across time as all EVs maintain their speeds. The power in the signal is concentrated at around the harmonics as predicted under scenario S4. Moreover, we still observe some spread around harmonic peaks due to EVs moving at time-varying speeds.
We also computed the spectrogram of the DWPT load to study how its frequency content varies with time. As pointed out in Fig. 8 under free-flowing traffic conditions, any $1$-min realization of the DWPT load would yield a similar spectrum and THC, as all EVs maintain their speeds across time.
@@
Congested traffic conditions. With the next test, we wanted to see if there exist traffic conditions under which the DWPT demand exhibits higher THC due to EV synchronization as in S1. To this end, we studied the DWPT load on an ER with heavy traffic moving at a slower speed. Such conditions can arise due to an accident, adverse weather conditions, or construction activities. To simulate such traffic conditions in SUMO, the speed limit was $24.56$ m/s ($55$ mph) on the first half of the ER and lowered on the second part. Initial EV speeds were drawn from the truncated Gaussian distribution explained earlier. The mean arrival rate was set to $\lambda_{a}=1$ EV/s. At this rate, the traffic jam propagated from the second half backwards to the beginning of the ER. Eventually, all EVs were moving at roughly the slower speed limit of the second half. Inter-vehicle spacings were found to be approximately constant across vehicles and time.
Figure 9: Variation of THC across time under low-speed, high-volume traffic simulated in SUMO. The speed affects the number of EVs and the EV spacing. Some speeds result in near-synchronization and high THC, as in S1, and other speeds result in lower THC, as in S4.
@@
As explained under S1, EV loads at constant speed can synchronize only when EV spacings are integer multiples of the coil segment length $D$. Moreover, the distance gap between EVs increases with the speed limit [16]. Based on this understanding, we identified several speed limits for the second half of the ER, at which approximate synchronization occurs and leads to high THC. Analogous to Fig. 8, Figure 9 depicts how the THC of the total DWPT load demand evolves across time for different jam speeds. The THC was computed based on $1$-min segments of $p(t)$ with a 50% overlap. We observe that higher speeds result in fewer EVs, as expected based on Remark 2. We also note that the THC scales as $\text{THC}_{h}/\sqrt{N}$ for the bottom three speed limits, for which the inter-vehicle distance was not a multiple of $D$. On the contrary, for the top three speed limits, THC increases due to the resonance of traffic with the ER infrastructure. The maximum THC value we observed was $18\%$ for a particular $1$-minute segment when the jam moves at $7.81$ m/s ($17.47$ mph). Such synchronization is impossible in free-flowing conditions, as EV spacings are randomized. Overall, this experiment validates the conjecture made in S1 that the DWPT load demand can exhibit high THC under heavy traffic conditions moving at specific speeds, especially at relatively lower speed limits. Also, the THC appears to be varying sinusoidally, which is a common phenomenon for frequency-modulated signals [26]. A detailed explanation of this sinusoidal variation in THC would be the topic of future study.
VIII Effects of DWPT Loads on a Power System
We simulated the impact of DWPT loads on the WECC 179-bus system using the ANDES grid dynamic simulator [9]. The WECC system has a base load of 60.78 GW and comprises 29 generators. To emulate the increasing penetration of inverter-based resources, the system’s inertia was reduced by 40%. The WECC system has unstable modes with negative damping ratios within the 0.1–0.7 Hz range [15]. In summary, our tests on the WECC indicate that, under congested traffic conditions, DWPT loads can excite grid frequency dynamics when their fundamental frequency is close to a critical mode. Moreover, distributing DWPT loads across geographically distant buses can exacerbate frequency oscillations. DWPT loads during free-flowing conditions are unlikely to excite any dynamics. The tests are described in detail next.
TABLE II: Frequency Content of DWPT Loads Powered by WECC
Cases Load THC (%) Load fundamental frequency (Hz)
Case A (3, 4, 5) (5.38, 5.86, 4.40)
Case B (15, 12, 12) (1.50, 2.11, 2.56)
Case C (3, 2, 3.5) (1.75, 2.25, 3.1)
Case D (14, 12, 12) (0.69, 2.11, 2.56)
In all tests, we considered DWPT systems connected at three buses. Although the bus locations varied per test as specified, the DWPT spectrum parameters were selected as shown in Table II. Case A simulates DWPT loads under free-flowing traffic conditions with DC power being approximately 10% of the nominal load demand of the hosting bus. Cases B-D simulate DWPT loads due to congested traffic with DC power being approximately 20% of the nominal load demand of the hosting bus. In Case C, DWPT loads do not synchronize as EV spacings are not integer multiples of the slab length $D$. DWPT loads were generated using vehicle trajectories simulated by SUMO applied to the INDOT ER specifications. To ensure that the DWPT load does not alter the system’s initial operating point or modes, we subtracted the DC term of the DWPT load from the original non-DWPT load at the three hosting buses.
Figure 10: Generator frequencies when DWPT loads are simulated per case B at: Top panel: nearby buses (54,58,60); and Bottom panel:) dispersed buses (54,100,164). The oscillations have a larger magnitude in the bottom panel; a maximum peak-to-peak deviation of 0.02 Hz, compared with 0.01 Hz in the top panel. One plausible explanation is that geographically distributed pulsating loads are more likely to excite multiple modes, thereby exacerbating oscillations.
Effect of DWPT location. We first studied how the location of DWPT loads impacts grid dynamics by considering two sets of DWPT-serving buses: c1) buses (54,58,60) that are geographically concentrated; and c2) buses (54,100,164) that are geographically distributed. In either placement case, DWPT loads were simulated per Case B. Figure 10 shows generator frequencies in the respective cases. Although each DWPT load has the same DC power between the two cases, case c2) triggered stronger oscillations. In particular, we observed maximum peak-to-peak frequency deviation of 0.01 Hz in c1), whereas in c2), peak-to-peak oscillations of amplitude 0.02 Hz are observed. A similar behavior of geographically dispersed pulsating loads yielding stronger oscillations has been observed in the context of pulsating datacenter loads while training large-language models (LLMs) [15].
Figure 11: Generator frequencies when DWPT loads are simulated per Case A (top) and Case C (bottom) at buses (54,100,164). Although loads are distributed geographically, their harmonic content is low, resulting in lower oscillation amplitudes.
Effect of DWPT THC. We simulated DWPT loads of different THCs sited on the same set of buses (54,100,164), and studied their impact on grid frequency dynamics. DWPT loads were simulated per Cases A, B, and C. Comparing the bottom panel of Fig. 10 with Fig. 11, we observe that DWPT loads exhibiting higher THC (Case B) introduce sustained frequency oscillations of amplitude 0.08 Hz. On the other hand, the lack of EV synchronization in Cases A and C introduces milder harmonics, leading to decaying oscillations.
Figure 12: Generator frequencies when DWPT loads are generated per case D at buses (54,100,164). Because one of the DWPT loads has a fundamental frequency close to the unstable 0.66-Hz mode, the system becomes unstable.
Figure 13: DWPT loads generated per case D at buses (54,100,164) leading to inter-area oscillations. Heed generators (2,4) swing out of phase compared to generators (11,14). The two generator pairs are in different WECC zones.
Effect of DWPT fundamental frequency. We also simulated grid dynamics when the DWPT load spectra contained peaks close to modal frequencies. The reduced-inertia WECC system exhibits a critical mode at 0.66 Hz, with significant participation from generator 14, located near bus 54. At traffic congestion speeds of around 8 mph, we observed that the DWPT load possesses a fundamental frequency of 0.69 Hz along with a THC of 14%; see blue plot in Fig. 9. DWPT loads were generated per Case D at buses (54,100,160). The corresponding dynamics are shown on Fig. 12. The DWPT load excites an unstable mode, and the system becomes unstable with generator frequencies diverging. Furthermore, frequencies associated with generators (2,4) swing out of phase compared to generators (11,14) as depicted in Fig. 13. These sets of generators are in different WECC zones, indicating that the particular load configuration introduces inter-area oscillations.
IX Conclusions
@@
This work has developed statistical models for the total DWPT load to be consumed at the substation serving an ER. By postulating different distributions of EV timing, our analysis explains how the DWPT spectrum varies across different traffic conditions. The worst-case scenario for THC occurs when EVs move in a synchronized manner with specific inter-vehicle spacings. Simulations using SUMO demonstrate that near synchronization can occur under low-speed, high-volume traffic. When EVs move freely, the DWPT load exhibits a more favorable spectral profile, as confirmed by simulations of low-volume traffic in SUMO. The frequency content at non-DC frequencies scales with $\sqrt{N}$ or $N$, depending on EV synchronism. Platoon formations can yield higher THC. In general, the fundamental frequency depends on the spacing between coils and the average EV speed, so that low-speed traffic can result in harmonics below 2 Hz. Higher-order harmonics decrease in magnitude and spread across a wider bandwidth. Despite the simplicity of the models, our findings provide valuable insights to ER planners and grid operators on how the DWPT spectrum scales with the number of EVs. We also studied the impact of DWPT loads on the WECC system and identified that DWPT loads during congested traffic conditions could excite system dynamics. Geographic load spreading has been shown to worsen oscillations.
Some open research directions include: a) Model DWPT loads as non-stationary signals and analyze their frequency content using tools such as the wavelet transform; b) Study the effect of correlated EV speeds based on dynamical models of traffic flows; c) Analyze the effect of DWPT load consumed at multiple buses on grid frequency dynamics; d) Develop solutions to smooth out DWPT load; e) Explore randomized triggering of Tx coils to naturally smooth out DWPT loads; and f) develop solutions to actively monitor DWPT loads for efficient system operation.
X Appendix
@@
where $a_{i}$ is defined as the effective rating of platoon $i$. Based on this decomposition, the DWPT load is now amenable to the analysis of Lemma 2 subject to two modifications. First, signal $p(t)$ is now the sum of $N/Q$ rather than $N$ stochastic components, each with i.i.d. and uniformly distributed timings $t_{i}$. Second, the effective rating $a_{i}$ of each stochastic component has a mean $Q\mu_{a}$, and its variance is increased to $Q\sigma_{a}^{2}$. The mean and autocorrelation of $p(t)$ can be obtained as in Lemma 2 by simply replacing $N$ by $N/Q$, $\mu_{a}$ by $Q\mu_{a}$, and $\sigma_{a}^{2}$ by $Q\sigma_{a}^{2}$. The autocorrelation function is